Find an equation of the line
a. passing through (-5,4) and parallel to 2x-3y=4. leave answer in slope intercept form.
I have worked the entire problem out and came up with y=2/3x-29/3. But the answer on the handout say it should be y=2/3x-22/3. what am i doing wrong?

Answer by longjonsilver(2297)

looks like an error in the working out of the fractional value of c.

2x-3y=4
2x-4=3y
3y=2x-4
y=(2/3)x-(4/3)

Its gradient is 2/3. As our line is parallel, its gradient is also 2/3. So we have:

y=mx+c
y=(2/3)x+c

We also know that the line passes through (-5,4), so we then get:
4=(2/3)(-5)+c
4=(-10/3)+c
c = 4+(10/3)
c = (12/3)+(10/3)
c = 22/3

--> y=(2/3)x+(22/3)

jon